Determine the energy of a photon with a wavelength of 330 nm.
Answer #1

Given:

lambda = 330 nm = 3.30*10^-7 m

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(3.30*10^-7 m)

= 6.02*10^-19 J

Answer: 6.02*10^-19 J